find the scalar product Ā•B of the two vectors. The magnitudes of the vectors are A=4.00 and B=5.00
1
Expert's answer
2019-07-10T10:39:25-0400

The scalar product by definition

"({\vec A}\cdot{\vec B})=AB\cos\theta"

So

"({\vec A}\cdot{\vec B})=4.00\times 5.00\cos\theta"

"=20.0\cos\theta"

Here theta is an angle between vectors A and B.